# R-Zero: Self-Evolving Reasoning LLM from Zero Data

### v2

@@ 1 Introduction
Self-evolving Large Language Models (LLMs) represent a promising frontier for advancing language intelligence. By autonomously generating, refining, and learning from their own experiences, these models provide a scalable pathway toward artificial superintelligence (Tao et al., 2024; Tan et al., 2024a). A critical requirement for training such self-evolving LLMs is access to large volumes of expertly curated tasks and labels, which serve as supervision signals for fine-tuning or reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; DeepSeek-AI et al., 2025). However, relying on human annotators to create these tasks and labels is not only costly, labor-intensive, and difficult to scale, but also presents a fundamental bottleneck to advancing AI systems toward capabilities that could eventually surpass human intelligence (Su et al., 2025; Zhao et al., 2025a).
@@ 3.2 Challenger Training
This function incentivizes questions where the Solver is maximally uncertain (accuracy approaches 50%). We provide a theoretical motivation for this reward function in Sec. 3.5.
@@ 4.1.3 Training Details
Our entire framework is implemented based on the EasyR1 codebase (Zheng et al., 2025b). In each iteration of the R-Zero co-evolutionary loop, we follow a specific set of hyperparameters. The Challenger ($Q_{\theta}$) first generates a candidate pool of $N=8,000$ questions. To construct the training dataset for the Solver, these questions are filtered based on consistency. For each candidate question, we sample $m=10$ answers from the current Solver ($S_{\phi}$). A question is retained for the training set only if the number of answers matching the majority-vote pseudo-label is between 3 and 7, inclusive ($\delta=0.25$). This numerical range is consistent with the methodology used in previous research (Zhang & Zuo, 2025; Li et al., 2025b; Bercovich et al., 2025). When training the Challenger, the uncertainty reward $r(x;\phi)$ is calculated by sampling $m=10$ responses from the Solver. For the intra-batch repetition penalty, we set the clustering distance threshold to $\tau_{\text{BLEU}}=0.5$. Further implementation details and prompts can be found in Appendix A.
@@ 4.3 Results in General Reasoning
Table 2: Results on general-domain reasoning benchmarks. The table compares the Base Model, a Base Challenger baseline, and our iterative R-Zero. The peak performance achieved during each model’s training process is highlighted in bold.
Model Name Overall AVG MATH AVG SuperGPQA MMLU-Pro BBEH
@@
   Base Model 27.10 42.58 20.88 37.38 7.57
   Base Challenger 30.83 44.36 24.77 47.59 6.59
   R-Zero (Iter 1) 34.27 48.06 27.92 51.69 9.42
   R-Zero (Iter 2) 34.92 48.44 27.72 53.75 9.76
   R-Zero (Iter 3) 34.64 49.07 27.55 51.53 10.42
@@
   Base Model 34.49 49.18 28.33 51.80 8.63
   Base Challenger 36.43 51.87 30.12 54.14 9.60
   R-Zero (Iter 1) 37.93 53.39 31.26 57.17 9.91
   R-Zero (Iter 2) 38.45 53.84 31.58 58.20 10.20
   R-Zero (Iter 3) 38.73 54.69 31.38 58.23 10.60
@@
   Base Model 12.27 26.64 10.09 10.87 1.46
   Base Challenger 14.41 27.51 11.19 14.53 4.40
   R-Zero (Iter 1) 14.93 27.76 12.21 15.72 4.05
   R-Zero (Iter 2) 15.11 28.20 12.43 16.08 3.74
   R-Zero (Iter 3) 15.67 29.32 12.44 16.71 4.20
@@
   Base Model 16.81 32.11 13.26 20.21 1.64
   Base Challenger 25.08 36.41 16.99 41.46 5.46
   R-Zero (Iter 1) 26.44 37.80 19.15 42.05 6.77
   R-Zero (Iter 2) 26.77 38.23 19.27 41.34 8.25
   R-Zero (Iter 3) 26.88 38.52 19.82 40.92 8.25
@@ 5.4.2 Beyond Label Noise: Unpacking the Roots of Instability
This suggests that the absolute percentage of label noise is not the sole determinant of instability. Another potential, and perhaps more fundamental, reason is a form of model collapse that can be introduced when training exclusively on self-synthesized data (Tan et al., 2024b; Shumailov et al., 2024; Dohmatob et al., 2024b; Zhou et al., 2025b; Seddik et al., 2024; Dohmatob et al., 2024a; Briesch et al., 2023; Zheng et al., 2025a). A model can enter a degenerative feedback loop, suffering from a loss of diversity or an amplification of its own biases, which presents a significant challenge.
@@ 6.2 Self-Play in Large Language Models
The paradigm of self-play, where models take on dual roles to create a self-improvement loop, has recently been adapted to improve language models without human data. This approach has been particularly fruitful in verifiable domains like code generation, where a “Coder” agent’s program is verified by a “Tester” agent’s unit tests (Lin et al., 2025; Wang et al., 2025a; Pourcel et al., 2025). More advanced frameworks push autonomy further by learning to generate the problems themselves, creating an adaptive curriculum from a small seed of examples or from scratch (Zhao et al., 2025a; Li et al., 2025c; Zhou et al., 2025a; Fang et al., 2025). Our work distinguishes itself by extending this paradigm to general reasoning domains that lack such verifiable environments, instead learning from a reward signal derived from the model’s own internal consistency.

### v3

@@ 1 Introduction
Self-evolving Large Language Models (LLMs) represent a promising frontier for advancing language intelligence. By autonomously generating, refining, and learning from their own experiences, these models provide a scalable pathway toward artificial superintelligence (Tao et al., 2024; Tan et al., 2024b). A critical requirement for training such self-evolving LLMs is access to large volumes of expertly curated tasks and labels, which serve as supervision signals for fine-tuning or reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; DeepSeek-AI et al., 2025). However, relying on human annotators to create these tasks and labels is not only costly, labor-intensive, and difficult to scale, but also presents a fundamental bottleneck to advancing AI systems toward capabilities that could eventually surpass human intelligence (Su et al., 2025; Zhao et al., 2025a).
@@ 3.2 Challenger Training
This function incentivizes questions where the Solver is maximally uncertain (accuracy approaches 50%) (Sun et al., 2025). We provide a theoretical motivation for this reward function in Sec. 3.5.
@@ 4.1.3 Training Details
Our entire framework is implemented based on the EasyR1 codebase (Zheng et al., 2025b). In each iteration of the R-Zero co-evolutionary loop, we follow a specific set of hyperparameters. The Challenger ($Q_{\theta}$) first generates a candidate pool of $N=8,000$ questions. To construct the training dataset for the Solver, these questions are filtered based on consistency. For each candidate question, we sample $m=10$ answers from the current Solver ($S_{\phi}$). A question is retained for the training set only if the number of answers matching the majority-vote pseudo-label is between 3 and 7, inclusive ($\delta=0.25$). This numerical range is consistent with the methodology used in previous research (Zhang and Zuo, 2025; Li et al., 2025b; Bercovich et al., 2025). When training the Challenger, the uncertainty reward $r(x;\phi)$ is calculated by sampling $m=10$ responses from the Solver. For the intra-batch repetition penalty, we set the clustering distance threshold to $\tau_{\text{BLEU}}=0.5$. Further implementation details and prompts can be found in Appendix A.
@@ 4.3 Results in General Reasoning
Table 2: Results on general-domain reasoning benchmarks. The table compares the Base Model, a R-Zero ([Uncaptioned image] challenger) baseline, and our R-Zero. The peak performance is highlighted in bold.
Model Name Overall Avg. SuperGPQA MMLU-Pro BBEH
@@
   Base Model (w/o training) 26.34 20.88 50.58 7.57
   R-Zero ([Uncaptioned image] challenger) 28.52 24.77 54.20 6.59
    R-Zero (our method) 31.15 27.55 55.47 10.42
@@
   Base Model (w/o training) 31.98 28.33 58.97 8.63
   R-Zero ([Uncaptioned image] challenger) 31.29 30.12 54.14 9.60
    R-Zero (our method) 34.50 31.38 61.53 10.60
@@
   Base Model (w/o training) 7.47 10.09 10.87 1.46
   R-Zero ([Uncaptioned image] challenger) 10.04 11.19 14.53 4.40
    R-Zero (our method) 11.12 12.44 16.71 4.20
@@
   Base Model (w/o training) 11.70 13.26 20.21 1.64
   R-Zero ([Uncaptioned image] challenger) 21.30 16.99 41.46 5.46
    R-Zero (our method) 23.00 19.82 40.92 8.25
@@ 5.4.2 Beyond Label Noise: Unpacking the Roots of Instability
This suggests that the absolute percentage of label noise is not the sole determinant of instability. Another potential, and perhaps more fundamental, reason is a form of model collapse that can be introduced when training exclusively on self-synthesized data (Tan et al., 2024a; Shumailov et al., 2024; Dohmatob et al., 2024b; Zhou et al., 2025b; Seddik et al., 2024; Dohmatob et al., 2024a; Briesch et al., 2023; Zheng et al., 2025a). A model can enter a degenerative feedback loop, suffering from a loss of diversity or an amplification of its own biases, which presents a significant challenge.
@@ 6.2 Self-Play in Large Language Models
The paradigm of self-play, where models take on dual roles to create a self-improvement loop (Chen et al., 2024), has recently been adapted to improve language models without human data. This approach has been particularly fruitful in verifiable domains like code generation, where a “Coder” agent’s program is verified by a “Tester” agent’s unit tests (Lin et al., 2025; Wang et al., 2025a; Pourcel et al., 2025). More advanced frameworks push autonomy further by learning to generate the problems themselves, creating an adaptive curriculum from a small seed of examples or from scratch (Zhao et al., 2025a; Li et al., 2025c; Zhou et al., 2025a; Fang et al., 2025). Our work distinguishes itself by extending this paradigm to general reasoning domains that lack such verifiable environments, instead learning from a reward signal derived from the model’s own internal consistency.